Nabor Industries is considering going public but is unsure of a fair offering price for the company. Before hiring an investment banker to assist in making the public​ offering, managers at Nabor have decided to make their own estimate of the​ firm's common stock value. The​ firm's CFO has gathered data for performing the valuation using the free cash flow valuation model. The​ firm's weighted average cost of capital is 11 %​, and it has $ 2 comma 670 comma 000 of debt at market value and $ 530 comma 000 of preferred stock in terms of market value. The estimated free cash flows over the next 5​ years, 2020 through 2024​, are given in the​ table,

Year​(t​)

Free cash flow

​(FCF​)

2020 200,000

2021 240,000

2022 310,000

2023 370,000

2024 410,000

Beyond 2024 to​ infinity, the firm expects its free cash flow to grow by 5 % annually. a.  Estimate the value of Nabor​ Industries' entire company by using the free cash flow valuation model. b.  Use your finding in part a​, along with the data provided​ above, to find Nabor​ Industries' common stock value. c.  If the firm plans to issue 200 comma 000 shares of common​ stock, what is its estimated value per​ share?
Answer #1

1.
=200000/1.11+240000/1.11^2+310000/1.11^3+370000/1.11^4+410000/1.11^5+410000/1.11^5*1.05/(11%-5%)=5346697.676

2.
=5346697.676-2670000-530000=2146697.676

3.
=2146697.676/200000=10.73348838